table for the questions

(Trillions of Dollars)

Real

GDP

(Y)

Net

Taxes

(NT)

Dispos

-able

Income

(Y-NT)

Con-

sumption

(C)

Saving

(S)

Planned

Invest-

ment

(I)

Govern-

ment

Purchases

(G)

Net

Exports

(X-M)

Planned

Aggregate

Expenditures

(C+I+G+(X-M)

5.0

1.0

4.0

3.9

0.1

1.0

1.0

-0.7

5.2

5.5

1.0

4.5

4.3

0.2

1.0

1.0

-0.7

5.6

6.0

1.0

5.0

4.7

0.3

1.0

1.0

-0.7

6.0

6.5

1.0

5.5

5.1

0.4

1.0

1.0

-0.7

6.4

7.0

1.0

6.0

5.5

0.5

1.0

1.0

-0.7

6.8

   19.   In Exhibit 10-2, the equilibrium level of GDP is

a.

$5.0 trillion

b.

$5.5 trillion

c.

$6.0 trillion

d.

$6.5 trillion

e.

$7.0 trillion

   20.   At the equilibrium level of GDP in Exhibit 10-2, injections equal

a.

$1.0 trillion

b.

$1.5 trillion

c.

$1.4 trillion

d.

$1.2 trillion

e.

$1.3 trillion

   21.   At the equilibrium level of GDP in Exhibit 10-2, leakages equal

a.

$1.4 trillion

b.

$1.3 trillion

c.

$1.5 trillion

d.

$1.0 trillion

e.

$1.1 trillion

   22.   At the equilibrium level of GDP in Exhibit 10-2, saving equals

a.

$0.4 trillion

b.

$0.6 trillion

c.

$0.5 trillion

d.

$0.2 trillion

e.

$0.3 trillion

   23.   In equilibrium in Exhibit 10-2, S + NT = I + G + (X - M) and is equal to

a.

$1.1 trillion

b.

$1.2 trillion

c.

$1.3 trillion

d.

$1.4 trillion

e.

$1.5 trillion

   24.   The marginal propensity to consume (MPC) in Exhibit 10-2 equals

a.

0.20 or 1/5

b.

0.40 or 2/5

c.

0.80 or 4/5

d.

0.90 or 9/10

e.

0.60 or 3/5

   25.   The marginal propensity to save (MPS) in Exhibit 10-2 equals

a.

0.80 or 4/5

b.

0.60 or 3/5

c.

0.40 or 2/5

d.

0.10 or 1/10

e.

0.20 or 1/5

   26.   In Exhibit 10-2, which group of the following is considered autonomous with respect to income?

a.

C, I, (X - M)

b.

S, I, G, (X - M)

c.

NT, C, G, (X - M)

d.

NT, S, I, (X - M)

e.

I, G, X
Answer #1

19.   In Exhibit 10-2, the equilibrium level of GDP is $6.5 trillion (Planned Aggregate Expenditures (C+I+G+(X-M) = Real GDP

20. At the equilibrium level of GDP in Exhibit 10-2, injections equal $1.3 trillion ( Ig + G + Xn).

21. At the equilibrium level of GDP in Exhibit 10-2, leakages equal 1.3 trillion (S+NT).

22. At the equilibrium level of GDP in Exhibit 10-2, saving equals 0.3 trillion.

23. In equilibrium in Exhibit 10-2, S + NT = I + G + (X - M) and is equal to 1.3 trillion.

24. The marginal propensity to consume (MPC) in Exhibit 10-2 equals MPC = 0.4/0.5 = 0.8

24. The marginal propensity to save (MPS) in Exhibit 10-2 equals 1-MPC = 1-0.8 = 0.2